# Randomization-based Inference for MCP-Mod

### v1

@@ 2 A Phase II Dose Finding Trial
The trial is a randomized, double-blind investigation. In this paper we assume that the trial is designed to assess the efficacy and safety of a drug across three different dose levels: 10 mg, 25 mg, and 100 mg, compared to a placebo (0 mg). We have in total $n=49$ patients, which were randomized using a $1:2:2:2$ randomization ratio (placebo against the three active doses). Permuted block randomization with blocks of fixed length of $B=7$ was employed to sequentially perform treatment assignments and to ensure that 7 patients were allocated to the placebo group and 14 patients to each of the active doses. The primary endpoint of the trial is binary, where a value of $1$ indicates a successful treatment outcome and a value of 00 treatment failure. The trial design aims to achieve $80$% power to detect the presence of dose-response signal at a one-sided significance level of $10$% using the generalized MCP-Mod approach. The assumed response rates are $20$% for the placebo group and $80$% for the highest dose group (100 mg). These parameters provide the foundation for the simulation study in Section 4. In the next Section we discuss and introduce methodology to analyze this trial example.
@@ 3 Methodology
In this section, we briefly introduce the generalized MCP-Mod approach, describe the challenges in estimation with binary outcomes for small sample sizes, and explain the implementation of randomization tests. Throughout this section we consider a parallel group, randomized, placebo-controlled trial, where $n$ patients are randomized among doses $d_{0},\ldots,d_{k-1}$, with $d_{0}$ representing the placebo dose. The treatment assignment for patient $i$ is indicated by $Z_{i}$, which can take any value from the set $\{d_{0},\ldots,d_{k-1}\}$. Collectively, the treatment assignments are represented by the vector $\mathbf{Z}=(Z_{1},\ldots,Z_{n})$, which is drawn from the randomization probability distribution $\mathcal{P}(\mathbf{Z})$ defined over all possible treatment sequences. Each patient $i$ has $k$ potential outcomes $y_{i,z}$ corresponding to each treatment $z$ within $\{d_{0},\ldots,d_{k-1}\}$. Additionally, patient $i$ has a covariate $\mathbf{x}_{i}$. Upon realization of the treatment sequence $\mathbf{z}$, only the outcome $y_{i,z_{i}}$ is observed for each patient $i$ and assigned treatment $z_{i}$; all other potential outcomes for that patient remain unobserved.
@@ 3.2 Challenges in Estimation with Binary Outcomes for Small Samples
In this section, we discuss complete or quasicomplete separation in logistic regression, which leads to non-existence of unique MLEs [17]. Complete separation in binary classification models refers to a situation where the predictor variables perfectly predict the outcome variable. In other words, there exists a hyperplane that can separate the two classes without any errors. Meaning that if there is a combination of predictor variables such that all observations with value $1$ lie on one side of the hyperplane and all observations with value 00 lie on the other side, then complete separation is achieved. Quasicomplete separation is a related scenario where some data points may lie exactly on the hyperplane, which also leads to non-existence of MLEs.
Heinze et al. (2020) [18] propose using Firth’s method [19], which modifies the score function of logistic regression to reduce the bias of MLEs, which tend to be biased away from zero. This approach ensures finite estimates closer to zero. Specifically, estimates using Firth’s method always exist, even in situation of complete or quasicomplete separation. The estimates can be calculated efficiently using the logistf-package in R [20].
@@ 3.3.2 Randomization-based inference for MCP-Mod
Let $\mathbf{D}_{\mathbf{Z}}$ be the $(n\times k)$ matrix that has entries $d_{i,j}=1$ if $Z_{i}=d_{j}$ and 00 otherwise. Furthermore, let $\mathbf{d}(\mathbf{Z})_{i}=(d_{i,0},\ldots,d_{i,k-1})$. A generalized linear model with linear predictor $\eta_{i}=\mathbf{d}(\mathbf{Z})_{i}^{\prime}\boldsymbol{\delta}+\mathbf{x}_{i}^{\prime}\boldsymbol{\beta}$ is fitted, where $\boldsymbol{\delta}=(\delta_{1},...,\delta_{k-1})^{\prime}$ contains the treatment differences versus placebo. Based on fitting this model (using either likelihood or penalized likelihood estimation), population-average model predictions can be performed based on the obtained parameter estimates. That is, the population average estimate for dose $j$ on the linear predictor scale would be $\delta_{j}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{x}_{i}^{\prime}\boldsymbol{\beta}$ and $\frac{1}{n}\sum_{i=1}^{n}\mathbf{x}_{i}^{\prime}\boldsymbol{\beta}$ for the placebo group, providing an estimate of the group means $\widehat{\boldsymbol{\mu}}_{\mathbf{Z}}$ with covariance matrix $\mathbf{S}_{\mathbf{Z}}$ calculated under the assumptions of the used model. The test statistic is then given by
@@ 3.3.3 Randomization Procedures
The other two procedures we consider in this paper address this issue by enforcing a pre-specified final allocation. One such procedure, the random allocation rule (RA), can be formalized as an urn design. In this design, the number of balls for each treatment corresponds to the number of patients intended for that treatment. We then would repeatedly draw balls at random, without replacement until the urn is empty, ensuring each treatment group receives its target quota of patients. The number of different allocation sequences for the RA procedure is $|\mathcal{R}_{Z}|=\binom{n}{n_{0}\;n_{1}\;\cdots\;n_{k-1}}=\frac{n!}{n_{0}!n_{1}!\cdots n_{k-1}!}$, where $n_{j}\geq 1$ is the final target sample size for the $j$-th group, $\sum_{j=0}^{k}n_{j}=n$, and each sequence in the reference set is equally likely to occur, with probability $\binom{n}{n_{0}\;n_{1}\;\cdots\;n_{k-1}}^{-1}$. While the RA procedure achieves exactly the target allocation numbers for the treatment groups, it may still result in non-negligible imbalances between treatments during the trial, which can pose challenges when adjusting for time trends. However, this procedure has several disadvantages. For one, it can lead to selection bias, as the allocation of patient $i$ is not independent of previous allocations. In extreme cases, when the urn contains balls of only one treatment, this can result in guaranteed predictions. Additionally, this procedure can still result in significant imbalances between treatments during the trial, which can pose challenges when adjusting for time trends.
The third procedure, the permuted block design (PBD), addresses this issue by repeatedly executing the random allocation rule with smaller blocks or urns. This procedure involves selecting a block size $B$ such that the total number of patients $n$ can be divided into $n/B$ identical blocks. These smaller blocks mirror the larger urn used in RA in terms of the proportion of patients allocated to each treatment. By carefully choosing a sufficiently small block size, this method helps prevent substantial imbalances throughout the trial. However, it also increases the predictability of the allocation of the next patient towards the end of each block, thereby raising the risk of selection bias. The number of different randomization sequences for a PBD with $\ell$ blocks of length $m$ (such that $\ell m=n$ and each block has the allocation ratio $m_{0}:m_{1}:\cdots:m_{k}$, with $\sum_{j=0}^{k}m_{j}=m$) is $\binom{m}{m_{0}\;m_{1}\;\cdots\;m_{k-1}}^{\ell}$ and all sequences are equally likely. Due to those advantages, PBD is the most used design in practice [29].
@@ 4.1 Simulations for Binary Endpoint
In this section, we use the clinical trial example from Section 2 and consider in total 14 scenarios to investigate the properties of randomization-based tests for MCP-Mod, see Table 1. We investigate three sample size scenarios. Doubling the original trial size provides insights into a moderate sample, while increasing the trial size tenfold allows us to examine asymptotic properties. This approach is driven by the notably small sample size of the original trial, which is why we do not consider exploring smaller trial sizes. The simulation scenarios are designed to achieve a similar asymptotic power for the larger trials sizes with success probabilities on the highest dose of $p_{k-1}=0.61$ ($n=98$) and $p_{k-1}=0.364$ ($n=490$) compared to $p_{k-1}=0.8$ for $n=49$, which is used to compare the simulated power in Table 2. In addition to PBD, we examine the performance of RA for all and CR for the largest trial sizes only. For the two smaller trial sizes, the likelihood of allocating very few patients to the control arm is too high when using CR. In the case of the smallest sample size, there is even a possibility that only one or none of the patients would be allocated to the control arm. Lastly, we also rerun all the scenarios with the inclusion of a time trend.
@@
Figure 1: Five candidate models for MCP-Mod.
In each scenario of the 14 scenarios, outcomes are simulated using an Emax model with $\theta_{2}=10$. We have $\theta_{0}=\log\frac{p_{0}}{1-p_{0}}$ and then determine $\theta_{1}$ by solving
$\log\frac{p_{k-1}}{1-p_{k-1}}=\log\frac{p_{0}}{1-p_{0}}+\theta_{1}\cdot\frac{d_{k-1}}{10+d_{k-1}}.$ (6)
We then plug the vector of treatment sequences $Z=(Z_{1},...,Z_{n})$ into the Emax function in order to derive individual success probabilities. We also examine the influence of one covariate. Covariates can be utilized to incorporate patient characteristics, including demographic factors such as age, gender, and race. Baseline measurements pertaining to the disease in question, such as weight, blood pressure or race can also be included [30]. In addition biomarkers can be added as important covariates [31]. In this simulation study, we draw values $x_{1},...x_{n}$ from a standard normal distribution and subsequently calculate the values $y_{i}=f(Z_{i},\theta)+0.6x_{i}$ for $i=1,..,n$. Applying the inverse of the logit function to these values provides the individual probabilities of success $\gamma_{1},...,\gamma_{n}$. The factor of $0.6$ is selected so that the covariate has a predictive strength of $AUC=0.66$ for the control group, which is motivated by a real trial example [31].
@@
In conclusion, our evaluation highlighted that residual-based randomization tests offer superior power and robust type-I error control compared to the population-based MCP-Mod method. These tests effectively mitigate issues of complete separation when penalized MLE are used and are computationally efficient, making them well-suited for clinical trials with varied conditions and sample sizes. Additionally, while the presence of time trends and different randomization methods can affect performance, residual-based tests consistently demonstrated reliability and effectiveness across various scenarios.
@@ 4.2 Pharmacometric Simulation Study
In this section, we explore the potential of randomization tests for MCP-Mod in a simulation scenario motivated by a pharmacometric model with a continuous endpoint. We conduct a simulation study inspired by the example presented by Buatois et al. (2021) [12], who proposed cLRT-Mod, a pharmacometric adaptation of the MCP-Mod methodology. The simulation in this section does not use a population model to simulate new outcome data for each simulated trial, but rather generates all potential outcomes for every patient once based on a pharmacometric model using the quasi-Monte Carlo method. Then each simulated trial only differs by the generated treatment sequence, i.e which potential outcome is observed for each patient. To our knowledge, the synergies between pharmacometric modelling and randomization tests have not yet been explored. Pharmacometric modelling often focuses on modelling individual dose-response curve (via use of covariates and random effects), this makes it ideally suited to generate potential outcomes under all doses for an individual patient.
We base our simulation study on the hypothetical phase II dose-finding trial for the development of a new monoclonal antibody for the treatment of wet-AMD [34] presented by Buatois et al. (2021) [12]. This disorder is characterized by the growth of abnormal blood vessels beneath the retina, damaging the macula and impairing visual acuity (VA). The aim of the new treatment is to target vascular endothelial growth factor (VEGF) in the retina to improve VA. The VA improvement is measured using the early treatment diabetic retinopathy study (ETDRS) chart, with scores based on the number of readable letters out of 70. The trial duration is set to 12 months with five parallel arms (placebo, or one of the following doses: 100, 200, 400, and 1000 $\mu$g). Observations are taken at baseline, day 7, and every month for 12 months. The baseline observation is included as a covariate in all data sets in this simulation study.
For our first data-generating scenario, we use the case of a few patients ($n=50$) and a strong drug effect [12]. The set of candidate models includes flat, linear, log-linear, Emax, and Sigmoid functions. The parameters for these models can be found in Buatois et al. (2021) [12]. The individual dose-response curves and thus potential outcomes for each patient generated from the pharmacometric model are shown in Figure 4, the average dose-response curve is shown in red.
Figure 4: Individual dose-response curves for the 50 patients. The average of the individual curves is shown in red.
@@
Table 3: For the randomization tests $n_{rand}=1,000$ re-randomizations are performed and the scenario was simulated $n_{sim}=10,000$ times. The Monte-Carlo error for power and type-I error is below 0.4% [33].
@@
In the second scenario presented in Table 3, we sort the potential outcomes based on the baseline observation, which can be interpreted as a positive time trend. One potential interpretation is that patients who join the trial later tend to have higher baseline observations, are in better health. This could be because patients whose visual acuity is most severely impacted by wet-AMD might enroll in the trial sooner. These patients might either volunteer earlier due to the severity of their condition or be prioritized for selection in hospitals, thus causing those with better visual acuity to enroll later.

### v2

@@ 2 A Phase II Dose Finding Trial
The trial is a randomized, double-blind investigation. In this paper we assume that the trial is designed to assess the efficacy and safety of a drug across three different dose levels: 10 mg, 25 mg, and 100 mg, compared to a placebo (0 mg). We have in total $n=49$ patients, which were randomized using a $1:2:2:2$ randomization ratio (placebo against the three active doses). For simplicity and the purpose here we assume permuted block randomization with blocks of fixed length of $B=7$ would be employed to sequentially perform treatment assignments and to ensure that 7 patients were allocated to the placebo group and 14 patients to each of the active doses. The primary endpoint of the trial is binary, where a value of $1$ indicates a successful treatment outcome and a value of 00 treatment failure. The trial design aims to achieve $80$% power to detect the presence of dose-response signal at a one-sided significance level of $10$% using the generalized MCP-Mod approach. The assumed response rates are $20$% for the placebo group and $80$% for the highest dose group (100 mg). These parameters provide the foundation for the simulation study in Section 4. In the next Section we discuss and introduce methodology to analyze this trial example.
@@ 3 Methodology
In this section, we briefly introduce the generalized MCP-Mod approach and address the challenges of estimation with binary outcomes in small sample sizes. To overcome these challenges, we adapt the MCP-Mod framework by incorporating two key methods: Firth’s method to handle complete separation and randomization-based inference to ensure a robust and valid inference framework. The latter provides appropriate type-I error rate control and enhances robustness against time trends, making it well-suited for small-sample settings.
Throughout this section we consider a parallel group, randomized, placebo-controlled trial, where $n$ patients are randomized among doses $d_{0},\ldots,d_{k-1}$, with $d_{0}$ representing the placebo dose. The treatment assignment for patient $i$ is indicated by $Z_{i}$, which can take any value from the set $\{d_{0},\ldots,d_{k-1}\}$. Collectively, the treatment assignments are represented by the vector $\mathbf{Z}=(Z_{1},\ldots,Z_{n})$, which is drawn from the randomization probability distribution $\mathcal{P}(\mathbf{Z})$ defined over all possible treatment sequences. Each patient $i$ has $k$ potential outcomes $y_{i,z}$ corresponding to each treatment $z$ within $\{d_{0},\ldots,d_{k-1}\}$. Additionally, patient $i$ has a covariate $\mathbf{x}_{i}$. Upon realization of the treatment sequence $\mathbf{z}$, only the outcome $y_{i,z_{i}}$ is observed for each patient $i$ and assigned treatment $z_{i}$; all other potential outcomes for that patient remain unobserved.
@@ 3.2 Challenges in Estimation with Binary Outcomes for Small Samples
In this section, we discuss complete or quasicomplete separation in logistic regression, which leads to non-existence of unique MLEs [17, 18]. Complete separation in binary classification models refers to a situation where the predictor variables perfectly predict the outcome variable. In other words, there exists a hyperplane that can separate the two classes without any errors. Meaning that if there is a combination of predictor variables such that all observations with value $1$ lie on one side of the hyperplane and all observations with value 00 lie on the other side, then complete separation is achieved. Quasicomplete separation is a related scenario where some data points may lie exactly on the hyperplane, which also leads to non-existence of MLEs.
Heinze et al. (2020) [19] propose using Firth’s method [20], which modifies the score function of logistic regression to reduce the bias of MLEs, which tend to be biased away from zero. This approach ensures finite estimates closer to zero. Specifically, estimates using Firth’s method always exist, even in situation of complete or quasicomplete separation. The estimates can be calculated efficiently using the logistf-package in R [21] or PROC LOGISTIC in SAS [22].
@@ 3.3.2 Randomization-based inference for MCP-Mod
Let $\mathbf{D}_{\mathbf{Z}}$ be the $(n\times k)$ matrix that has entries $d_{i,j}=1$ if $Z_{i}=d_{j}$ and 00 otherwise. Furthermore, let $\mathbf{d}(\mathbf{Z})_{i}=(d_{i,0},\ldots,d_{i,k-1})$. A generalized linear model with linear predictor $\eta_{i}=\mathbf{d}(\mathbf{Z})_{i}^{\prime}\boldsymbol{\delta}+\mathbf{x}_{i}^{\prime}\boldsymbol{\beta}$ is fitted, where $\boldsymbol{\delta}=(\delta_{0},...,\delta_{k-1})^{\prime}$ are the dose specific intercepts. Based on fitting this model (using either likelihood or penalized likelihood estimation), population-average model predictions can be performed based on the obtained parameter estimates. That is, the population average estimate for dose $j$ on the linear predictor scale would be $\delta_{j}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{x}_{i}^{\prime}\boldsymbol{\beta}$, providing an estimate of the group means $\widehat{\boldsymbol{\mu}}_{\mathbf{Z}}$ with covariance matrix $\mathbf{S}_{\mathbf{Z}}$ calculated under the assumptions of the used model. The test statistic is then given by
@@ 3.3.3 Randomization Procedures
The other two procedures we consider in this paper address this issue by enforcing a pre-specified final allocation. One such procedure, the random allocation rule (RA), can be formalized as an urn design. In this design, the number of balls for each treatment corresponds to the number of patients intended for that treatment. We then would repeatedly draw balls at random, without replacement until the urn is empty, ensuring each treatment group receives its target quota of patients. The number of different allocation sequences for the RA procedure is $|\mathcal{R}_{Z}|=\binom{n}{n_{0}\;n_{1}\;\cdots\;n_{k-1}}=\frac{n!}{n_{0}!n_{1}!\cdots n_{k-1}!}$, where $n_{j}\geq 1$ is the final target sample size for the $j$-th group, $\sum_{j=0}^{k}n_{j}=n$, and each sequence in the reference set is equally likely to occur, with probability $\binom{n}{n_{0}\;n_{1}\;\cdots\;n_{k-1}}^{-1}$. While the RA procedure achieves exactly the target allocation numbers for the treatment groups, it may still result in non-negligible imbalances between treatments during the trial, which can pose challenges when adjusting for time trends. However, this procedure has several disadvantages. For one, it can lead to selection bias, as the allocation of patient $i$ is not independent of previous allocations. In extreme cases, when the remaining balls in the urn correspond to only one treatment, this can result in guaranteed predictions.
The third procedure, the permuted block design (PBD), addresses this issue by repeatedly executing the random allocation rule with smaller blocks or urns. This procedure involves selecting a block size $B$ such that the total number of patients $n$ can be divided into $n/B$ identical blocks. These smaller blocks mirror the larger urn used in RA in terms of the proportion of patients allocated to each treatment. By carefully choosing a sufficiently small block size, this method helps prevent substantial imbalances throughout the trial. Due to this advantage, PBD is the most used design in practice [31]. However, it also increases the predictability of the allocation of the next patient towards the end of each block, thereby raising the risk of selection bias when the block size is too small. The number of different randomization sequences for a PBD with $\ell$ blocks of length $m$ (such that $\ell m=n$ and each block has the allocation ratio $m_{0}:m_{1}:\cdots:m_{k}$, with $\sum_{j=0}^{k}m_{j}=m$) is $\binom{m}{m_{0}\;m_{1}\;\cdots\;m_{k-1}}^{\ell}$ and all sequences are equally likely.
@@ 4.1 Simulations for Binary Endpoint
In this section, we use the clinical trial example from Section 2 and consider in total 14 scenarios to investigate the properties of randomization-based tests for MCP-Mod, see Table 1. We investigate three sample size scenarios. Doubling the original trial size provides insights into a moderate sample, while increasing the trial size tenfold allows us to examine asymptotic properties. This approach is driven by the notably small sample size of the original trial, which is why we do not consider exploring smaller trial sizes. For comparability across scenarios the response probability at the highest dose $p_{k-1}$ was changed for each sample size to achieve the same (asymptotic) power, for population based MCP-Mod, as for the scenario with the smallest trial size $(n=49)$, which used $p_{k-1}=0.8$. Calculations were performed using the MCP-Mod design application [32], resulting in success probabilities of $p_{k-1}=0.61$ for $n=98$ and $p_{k-1}=0.364$ for $n=490$. Differences in the reported power results in Table 2 arise due to complete separation issues and the distinction between finite-sample and asymptotic properties. In addition to PBD, we examine the performance of RA for all and CR for the largest trial sizes only. For the two smaller trial sizes, the likelihood of allocating very few patients to the control arm is too high when using CR. In the case of the smallest sample size, there is even a possibility that only one or none of the patients would be allocated to the control arm. Lastly, we also rerun all the scenarios with the inclusion of a time trend.
@@
Figure 1: Five candidate models for MCP-Mod (transformed from a logit scale to a probability scale for easier interpretability).
We choose the Emax model with $\theta_{2}=10$ as the ground truth for the 14 simulation scenarios to provide a common basis for comparing the population-based generalized MCP-Mod approach with our proposed randomization-based inference extensions. Since all methods utilize the same candidate models, the observed differences in power should not be driven by the choice of the underlying truth but rather by the differences between the testing approaches. For the Emax model with $\theta_{2}=10$, we obtain $\theta_{0}=\log(p_{0}/(1-p_{0}))$ and determine $\theta_{1}$ by solving
$\log\frac{p_{k-1}}{1-p_{k-1}}=\log\frac{p_{0}}{1-p_{0}}+\theta_{1}\cdot\frac{d_{k-1}}{10+d_{k-1}},$ (6)
e.g. $\theta_{0}\approx-1.39$, $\theta_{1}\approx 3.05$ and $\theta_{2}=10$ for the clinical trial example with $n=49$ and $p_{k-1}=0.8$. We then plug the vector of treatment sequences $Z=(Z_{1},...,Z_{n})$ into the Emax function in order to derive individual success probabilities. We also examine the influence of one covariate. Multiple covariates can be utilized to incorporate patient characteristics, including demographic factors such as age, gender, and race. Baseline measurements pertaining to the disease in question, such as weight, blood pressure or race can also be included [33]. In addition biomarkers can be added as important covariates [34]. In all scenarios of this simulation study we include just one covariate. For this, we draw values $x_{1},...x_{n}$ from a standard normal distribution and subsequently calculate the values $y_{i}=f(Z_{i},\theta)+0.6x_{i}$ for $i=1,..,n$. Applying the inverse of the logit function to these values provides the individual probabilities of success $\gamma_{1},...,\gamma_{n}$. The factor of $0.6$ is selected so that the covariate has a predictive strength of $AUC=0.66$ for the control group, which is motivated by a real trial example [34].
@@
In conclusion, our evaluation highlighted that residual-based randomization tests offer superior power and robust type-I error rate control compared to the population-based MCP-Mod method. These tests effectively mitigate issues of complete separation when penalized MLE are used, allow the inclusion of multiple covariates and are computationally efficient, making them well-suited for clinical trials with varied conditions and sample sizes. Additionally, while the presence of time trends and different randomization methods can affect performance, residual-based tests consistently demonstrated reliability and effectiveness across various scenarios. Although we also examined a population-based MCP-Mod approach using penalized MLEs during our investigation, our primary focus in this paper remained on the randomization-based framework. While the population-based MCP-Mod approach using penalized MLEs helps mitigate complete separation and offers some power gains, it remains less effective than randomization-based methods. Notably, the randomization-based inference framework ensures type-I error rate control, even in the presence of time trends, and generally exhibits greater robustness across various scenarios.
@@ 4.2 Pharmacometric Simulation Study
In this section, we explore the potential of randomization tests for MCP-Mod in a simulation scenario motivated by a pharmacometric model with a continuous endpoint. We conduct a simulation study inspired by the example presented by Buatois et al. (2021) [12], who proposed cLRT-Mod, a pharmacometric adaptation of the MCP-Mod methodology. The simulation in this section does not use a population model to simulate new outcome data for each simulated trial, but rather generates all potential outcomes for every patient once based on a pharmacometric model using a quasi-Monte Carlo method. Then each simulated trial only differs by the generated treatment sequence, i.e which potential outcome is observed for each patient. To our knowledge, the synergies between pharmacometric modeling and randomization tests have not yet been explored. Pharmacometric modeling often focuses on modeling the individual dose-response curves (via use of covariates and random effects), this makes it ideally suited to generate potential outcomes under all doses for an individual patient.
We base our simulation study on the hypothetical phase II dose-finding trial for the development of a new monoclonal antibody for the treatment of wet-AMD [37] presented by Buatois et al. (2021) [12]. This disorder is characterized by the growth of abnormal blood vessels beneath the retina, damaging the macula and impairing visual acuity (VA). The aim of the new treatment is to target vascular endothelial growth factor (VEGF) in the retina to improve VA. The VA improvement is measured using the early treatment diabetic retinopathy study (ETDRS) chart, with scores based on the number of readable letters out of 70. The trial duration is set to 12 months with five parallel arms (placebo, or one of the following doses: 100, 200, 400, and 1000 $\mu$g). Observations are collected at baseline, day 7, and monthly for 12 months. In our simulation study, the baseline observation is included as a covariate, and the month 12 observation as primary endpoint. All other monthly observations are not considered in the analysis.
For our first data-generating scenario, we use the case of a few patients ($n=50$) and a strong drug effect [12]. The set of candidate models includes flat, linear, log-linear, Emax, and Sigmoid functions. The parameters for these models can be found in Buatois et al. (2021) [12]. We generated all potential outcomes under the Emax model with $\theta_{2}=150$ and the individual dose-response curves and thus potential outcomes for each patient generated from the pharmacometric model are shown in Figure 4, the average dose-response curve is shown in red.
Figure 4: Individual dose-response curves for the 50 patients. The average of the individual curves is shown in red and the grey lines represent the dose levels that are included in the trial.
@@
Table 3: For the randomization tests $n_{rand}=1,000$ re-randomizations are performed and the scenario was simulated $n_{sim}=10,000$ times. The Monte-Carlo standard error for power and type-I error is below 0.4% [36].
@@
In the second scenario presented in Table 3, we sort the potential outcomes based on the baseline observation, which can be interpreted as a positive time trend. One potential interpretation is that patients who join the trial later tend to have higher baseline observations, and are in better health. This could be because patients whose visual acuity is most severely impacted by wet-AMD might enroll in the trial sooner. These patients might either volunteer earlier due to the severity of their condition or be prioritized for selection in hospitals, thus causing those with better visual acuity to enroll later.